OpenStudy (anonymous):

having a hard time with -2(a+5)=27-2a answer won't check

OpenStudy (anonymous):

FALSE

OpenStudy (anonymous):

I need to solve and came up with -4a=22, but not checking

OpenStudy (katrinakaif):

-2a -10 = 27 - 2a +0a = 37 a = 0 Its false or a is simply = 0